Find the magnitude of D=7i-3j-2k,, E=4i+5j-3k


1
Expert's answer
2021-07-14T10:16:00-0400

By definition, the magnitude of any vector is given as follows:


"|A| = \sqrt{a_x^2 + a_y^2 + a_z^2}"

where "a_x,a_y, a_z" are corresponding Cartesian components of the vector.

Thus, obtain:


"|D| = \sqrt{7^2 + (-3)^2 + (-2)^2}\approx 7.9\\
|E| = \sqrt{4^2 + 5^2 + (-3)^2}\approx 7.1"

Answer. 7.9 and 7.1.